remi2019:

What is the value of a in the equation a = 2 + 3a + 10?

jhonyy9:

like a first step add the 10 and 2

jhonyy9:

10+2 = ?

supie:

you would change the equation to \(a=(3a)+(2+10)\) Then combine like terms which is adding 2 and 10.

remi2019:

12

supie:

yes so \(a=3a+12\) so do you know the second step?

remi2019:

15a?

supie:

? no

remi2019:

i am confused

supie:

the next step is to subtract \(3a\) from both sides.

jhonyy9:

@supie i say subtract just an a from both sides and after this subtract 12 from both sides to get an a positive on right side

supie:

"i say subtract just an a from both sides"?

hhanan:

the answer is -6

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie "i say subtract just an a from both sides"? \(\color{#0cbb34}{\text{End of Quote}}\) one ,,a" from both sides and not 3a

supie:

It says `3a` tho

supie:

a = 2 + `3a` + 10

Timmyspu:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie the next step is to subtract \(3a\) from both sides. \(\color{#0cbb34}{\text{End of Quote}}\) Supie is right because if you get only one side of the equation then you will get a wrong answer.

supie:

yes @jhonyy9 ?

hhanan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan the answer is -6 \(\color{#0cbb34}{\text{End of Quote}}\)

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan \(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan the answer is -6 \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{End of Quote}}\) stop giving direct answers

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie a = 2 + `3a` + 10 \(\color{#0cbb34}{\text{End of Quote}}\) a-a = 12 +3a-a 0 = 12 +2a -12 = 12 -12 +2a -12 = 2a -12/2 = a a = -6 these is what i ve said above

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @supie "i say subtract just an a from both sides"? \(\color{#0cbb34}{\text{End of Quote}}\) one ,,a" from both sides and not 3a \(\color{#0cbb34}{\text{End of Quote}}\) you said `not 3a`

jhonyy9:

yes just one ,,a" like how i ve wrote above

supie:

I did 3a and I got the same answer ....